OpenStudy (anonymous):

Consider the function f(x)=(8-x)/(x+9)+2 Algebraically determine the domain and all of the intercepts.

hero (hero):

\[x + 9 \ne 0\]\[f(x) \ne -1\]

hero (hero):

To find the intercepts, set f(x) = 0, then solve for x. Then set x = 0, then solve for f(x)

OpenStudy (anonymous):

How did you get f(x)=-1?

hero (hero):

Just by looking at the function really. I know that x + 9 is in the denominator, but when you are solving for x you usually multiply both sides by x + 9 to get rid of the denominator. But if both sides of the equation have -x, then there is no solution. The only way to get -x on both sides is if f(x) = -1

hero (hero):

Actually, I didn't see the + 2

hero (hero):

That changes things

OpenStudy (anonymous):

Alrighty, well I go a different answer. When I solved for x I got -26 and for the domain I got x can't equal -9 so the domain would be [(-\infty,-9)(-9,\infty)\]

OpenStudy (anonymous):

opps \[(-\infty,-9)(-9,\infty)\]

hero (hero):

You found what x cannot equal, not what f(x) cannot equal.

hero (hero):

I agree that \(x \ne -9\)

hero (hero):

I don't think you have to worry about what \(f(x)\) cannot equal since it doesn't ask you to find the range.

hero (hero):

Your domain appears to be correct. Now if you manage to find the intercepts, then you'll be all set.

hero (hero):

Because of the + 2, \(f(x) \ne 1\). You would only need to know that if you were trying to find the range.

OpenStudy (anonymous):

Mmkay! Thanks. Now, I just gotta find the intercepts!

hero (hero):

I already posted what to do to find the intercepts.

OpenStudy (anonymous):

I know I'm solving it haha.

hero (hero):

Okay, good luck.

OpenStudy (anonymous):

I got x-intercept (-26,0) Y- intercept (0, 26/9)

hero (hero):

That appears to be correct.

OpenStudy (anonymous):

Thanks for the help.